A rectangle is twice as long as it is wide. If it’s length and width are both decreased by 4 cm, it’s area is decreased by 164cm². Find it’s original dimensions.
Answer by Alan3354(69443)

A rectangle is twice as long as it is wide. If it’s [sic] length and width are both decreased by 4 cm, it’s [sic] area is decreased by 164cm². Find it’s [sic] original dimensions.
------------
Area = L*W
(L - 4)*(W - 4) = L*W - 164
L*W - 4W - 4L + 16 = L*W - 164
4L + 4W = 180
L + W = 45
-----------
Any combination of L & W will work, if both L & W > 4.
------------
PS it's = it is